May 27-30

– Five stores provided ~41.5 hours of luminosity
– NuMI off for target replacement
– Tevatron electron lens noise caused beam problems
– Tevatron sector B3 wet engine repairs halted, B3 cryo system operated via Jewel Thompson (JT) valve
– Store 8770 quenched due to abort kicker prefire
– Lightening strikes interrupted shot setup and disrupted other systems

*The integrated luminosity for the period from 5/23/11 to 5/30/11 was 58.41 inverse picobarns.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts